# Epistemic Diversity and Knowledge Collapse in Large Language Models

### v5

@@ Abstract
Large language models (LLMs) tend to generate lexically, semantically, and stylistically homogenous texts. This poses a risk of knowledge collapse, where homogenous LLMs mediate a shrinking in the range of accessible information over time. Existing works on homogenization are limited by a focus on closed-ended multiple-choice setups or fuzzy semantic features, and do not look at trends across time and cultural contexts. To overcome this, we present a new methodology to measure epistemic diversity, i.e., variation in real-world claims in LLM outputs, which we use to perform a broad empirical study of LLM knowledge collapse. We test 27 LLMs, 155 topics covering 12 countries, and 200 prompt variations sourced from real user chats. For the topics in our study, we show that while newer models tend to generate more diverse claims, nearly all models are less epistemically diverse than a basic web search. We find that model size has a negative impact on epistemic diversity, while retrieval-augmented generation (RAG) has a positive impact, though the improvement from RAG varies by the cultural context. Finally, compared to a traditional knowledge source (Wikipedia), we find that country-specific claims reflect the English language more than the local one, highlighting a gap in epistemic representation. [^1] Code and data: https://github.com/dwright37/llm-knowledge
@@ 1 Introduction
Figure 1: In this work, we measure epistemic diversity – via variability in claims about the world – for characterizing knowledge collapse in LLMs.
@@
At the same time, recent studies have noted that LLM outputs are homogeneous Sourati et al. (2025b). For example, LLMs reflect only a narrow range of writing and reasoning styles, use a limited vocabulary Sourati et al. (2025a) and convey only certain semantics Lee et al. (2025). This phenomenon, where models regress to a “central tendency”, also affects knowledge i.e., the pieces of information that models tend to generate Dutta and Chakraborty (2023). This could in turn limit the information available to the general public and subsequently dwindle our collective knowledge, a phenomenon broadly defined as knowledge collapse (Fig. 1). [^2] For an in depth definition and discussion, see Peterson (2025). Recent work  Peterson (2025) has begun to theorize this, but offers limited empirical data. To better characterize and understand this problem, we investigate, for the first time, whether LLMs themselves exhibit knowledge collapse.
@@
To do so, we perform an empirical study measuring epistemic diversity, defined as the diversity of claims about the world in a given corpus (e.g., a set of LLM responses). To measure this, we develop a new methodology intended to reveal the diversity of claims occurring in free-text LLM outputs. This involves (1) sampling outputs from LLMs using a set of 200 natural writing assistance prompts collected from Röttger et al. (2025), (2) partitioning the LLM responses into unique classes of semantically equivalent claims Farquhar et al. (2024), and (3) quantifying the diversity of the sample of claims with Hill-Shannon diversity, a widely used metric for measuring species diversity in ecology Roswell et al. (2021). From among the Llama, Gemma, Qwen, and OpenAI model families, we study 27 LLMs spanning multiple versions, sizes, and release dates, querying them for 155 topics.
@@
For the selected topics, while we find that epistemic diversity has increased since 2023 for three of four model families, it remains low for all models compared to a rudimentary web search. Similar to contemporary work Zhang et al. (2025), we find that model size has a statistically significant negative impact on epistemic diversity – smaller models generate more diverse knowledge than larger ones. In contrast, retrieval-augmented generation (RAG) has a statistically significant positive impact, highlighting the importance of RAG in preventing future knowledge collapse. For country-specific topics, we find that RAG has an uneven effect; certain countries (e.g., the USA) see more benefit due to a greater diversity in their RAG sources. Finally, compared to a traditional knowledge source (Wikipedia) in both English and local languages for country-specific topics, we find that the claims generated in our study reflect English language knowledge more than local language knowledge, highlighting a gap in epistemic representation.
@@ 2 Background
An increasing body of work shows that LLMs suffer from a lack of diversity along many dimensions; for a recent survey, see Sourati et al. 2025b. This lack of diversity includes lexical and stylistic Sourati et al. (2025a), semantic Lee et al. (2025), creative Xu et al. (2025), and perspective diversity Wright et al. (2024); Abdurahman et al. (2024); Zhang et al. (2025); Durmus et al. (2023); Röttger et al. (2024); Moore et al. (2024). Our work is most similar to perspective diversity, which has shown that LLM-generated views, opinions, and beliefs tend to reflect only a small subset of the world Durmus et al. (2023); Atari et al. ; Abdurahman et al. (2024); Alvero et al. (2024). Perspective diversity is usually measured with multiple-choice survey responses Durmus et al. (2023) or free-text responses partitioned into different classes Zhang et al. (2025). In our work we measure the epistemic diversity of models by partitioning the LLM output space into clusters of claims. This allows us to avoid shallow and/or fuzzy features such as semantic and lexical similarity, and improves upon unreliable multiple-choice setups Röttger et al. (2024).
@@
LLMs can exacerbate their own biases, leading to model collapse Shumailov et al. (2024). A growing concern among scholars is that such homogenization, combined with increased adoption of LLMs, will lead to epistemic problems at a societal level Zheng and Lee (2023); Messeri and Crockett (2024); Peterson (2025); Wagner and Jiang (2025); Qiu et al. (2025). Peterson (2025) defines “knowledge collapse” as LLMs facilitating a dwindling of knowledge into an increasingly narrow set of ideas. Knowledge collapse may affect existing knowledge sources such as Wikipedia Wagner and Jiang (2025), erase minoritized knowledge Zheng and Lee (2023), pollute scientific discoveries Messeri and Crockett (2024), hamper political discourse Coeckelbergh (2025), and limit ideation in writing Anderson et al. (2024). These concerns echo other contemporary epistemic issues such as “popularity bias” Ciampaglia et al. (2018), and “filter bubbles” Nguyen et al. (2014) occurring with the use of recommender systems. To assess the risk of knowledge collapse, our work offers a new methodology for measuring epistemic diversity and an empirical study of this phenomenon in LLMs.
@@ 3 Problem Setup and Notation
We measure the epistemic diversity of LLMs as the variation in the claims that the models make across different topics and prompts. To do so, we develop a methodology that is naturalistic and theoretically motivated. It is naturalistic in that it uses natural prompts to sample open-ended responses Wright et al. (2024); Moore et al. (2024); Röttger et al. (2025). It is theoretically motivated in that it adopts a statistically grounded measure of diversity used widely in ecology for measuring species diversity Peterson (2025); Roswell et al. (2021). We measure epistemic diversity after transforming open-ended LLM responses into distributions over meaning classes Farquhar et al. (2024).
To describe our approach (Fig. 1), we adopt the following notation. First, a corpus of text $P_{m{}_{t{}}}{}$ is elicited about a topic $t$ from a model mm{}. $P_{m{}_{t{}}}{}$ contains free text which can be decomposed into a list of $n$ claims $C_{m{}_{t}}$, which can be further partitioned into a set of unique meaning classes $\mathcal{X}_{m{}_{t}}$. A unique meaning class is a cluster where all claims within a given class mutually entail each other and do not mutually entail claims in other classes. Then, $x_{i}\in\mathcal{X}_{m_{t}}$ is defined as the empirical frequency of meaning class $i$ calculated from $C_{m{}_{t}}$, and $p_{i}$ is the probability of $i$, calculated as the relative frequency of $i$ in the sample, i.e., $\frac{x_{i}}{n}$.
Why not semantic similarity?
An alternative version of our setup would be to partition claims based solely on semantic similarity, as previous work has done Lee et al. (2025); Wright et al. (2024). However, sentence embeddings Reimers and Gurevych (2019) can only measure equivalence based on functionally similar phrases. For example, the phrases “Claude Shannon is the father of information theory” and “Claude Shannon is not the father of information theory” have extremely high semantic similarity (0.94) [^3] Huggingface ID: all-MiniLM-L6-v2, even though they present opposing claims. Similarly, “Claude Shannon is the father of quantum theory” is an entirely different claim from “Claude Shannon is the father of information theory,” but these two also have high semantic similarity (0.805). This setup would not reflect our definition of meaning classes.
@@ 4 Data Collection
    Cluster: Group the list of atomic claims into meaning classes $\mathcal{X}_{m{}_{t}}$.
@@ 4.1 Generation and Decomposition
We use the open-ended writing assistance prompt templates collected in Röttger et al. (2025). The original prompt templates are sourced from WildChat Zhao et al. (2024), a collection of natural conversations between users and ChatGPT. From the original 1,000 writing assistance prompts, we manually select a subset of 479 prompts that primarily focus on information seeking and informational writing. We manually filter out NSFW templates, creative writing templates such as “write a 50’s soviet style song about tt,” outlines such as “write an index for a book on tt,” or those explicitly asking for references. We then randomly sample 200 of these prompt templates for generation.
After generating responses ($P_{m{}_{t{}}}{}$), we decompose each response into a list of individual claims $C_{m{}_{t}}$. We do so using a strong open-weight LLM, [^4] Huggingface ID: meta-llama/Llama-3.1-70B-Instruct prompting the model to decompose non-overlapping chunks of three sentences at a time. This allows us balance claim recall while with ensuring that each input chunk is decontextualized.
@@
To evaluate the quality of decomposition, we develop three initial decomposition prompts (P1-P3) and have two independent annotators label (1) the quality of individual decomposed claims on a Likert scale from 1-5, and (2) how many claims in the original input chunk are missing from the list of decomposed claims. (On our scale, 1 means that the decomposed claim is not inferable from the original text, and 5 means that the claim is fully inferable and decontextualized.) The full annotation instructions appear in the Appendix § A.1. We label 100 input chunks for quality (1) and 108 instances for missing (2), all anonymized for the prompt type (P1-P3). Between the two annotators, we achieve a Kendall-Tau ($\tau$) correlation Kendall (1938) of 0.53 for (1) and 0.39 for (2), indicating moderate to strong agreement. [^5] We select $\tau$ because it is suited for ordinal data with ranking ties Kendall (1945). We break ties by having a third annotator label those instances where the original annotators disagreed.
We then set up an LLM-as-a-judge using G-Eval Liu et al. (2023), which achieves 0.6 Pearson correlation with the human labels for decomposition quality and 0.68 for the number of missing claims. We use ths to automatically label 6k instances across three decomposition prompt variants (P1-P3; see Appendix § A.2 for prompt text). The results are given in Table 1. We use P3 for the final decomposition in order to prioritize quality.
@@ 4.2 Clustering
Algorithm 1 outlines our approach for clustering the decomposed claims $C_{m{}_{t}}$ into meaning classes $\mathcal{X}_{m{}_{t}}$
@@
Following Farquhar et al. (2024), we bootstrap clusters based on mutual entailment using a strong pretrained model for natural language inference. [^6] Huggingface ID: microsoft/deberta-large-mnli This is done by assuming that when a given claim mutually entails at least one claim in an existing cluster, it also entails all other claims in the cluster. To reduce measuring mutual entailment across $n^{2}$ pairs for a set of $n$ claims, we only check entailment between the $N$ most similar claims (i.e., NnN*n comparisons), where similarity is measured using a strong S-BERT model. [^7] Huggingface ID: all-MiniLM-L6-v2 To mitigate potential drift in cluster cohesion (i.e., large clusters containing multiple meaning classes), we perform a post-processing step to break up large clusters using DBSCAN Ester et al. (1996), a related clustering approach Peterson (2025); Wright et al. (2024).
@@
Two independent annotators label: (1) the degree of cohesion among the clustered claims (Likert score between 1 and 5), and (2) whether semantically similar, non-clustered claims are grouped with a given cluster (binary yes/no). See Appendix § A.1 for full instructions. We label 100 instances of cohesion (1) and 360 instances of missing (2), achieving a Kendall-Tau correlation of 0.38 for (1) and 87% agreement accuracy for (2). We again break ties using a third annotator and set up an LLM-as-a-judge using G-Eval. We develop prompts that achieve 0.68 Pearson $r$ with the human labels for measuring cluster cohesion and 83.5 weighted F1 score for measuring binary missing claims. 15,000 samples of 3-5 sentences are then automatically labeled from clusters acquired using Algorithm 1, where semantic similarity for task (2) is measured using S-BERT. We acquire a cluster cohesion score of 4.08 and a missing rate of 13.79 (out of 100), indicating that the cluster cohesion is generally of high quality while missing only a relatively small proportion of singletons.

### v6

@@ Abstract
Large language models (LLMs) tend to generate homogenous texts, which may impact the diversity of knowledge generated across different outputs. Given their potential to replace existing forms of knowledge acquisition, this poses a risk of knowledge collapse, where homogenous LLMs may lead most people to be exposed to largely the same information, thus mediating a shrinking in the range of accessible information over time as underepresented knowledge is forgotten. To assess the risk of knowledge collapse with LLMs, we present a new methodology to measure epistemic diversity, i.e., variation in real-world claims in LLM outputs. We use this to perform a broad empirical study testing 27 LLMs, 155 topics covering 12 countries, and 200 prompt templates sourced from real user chats. For the topics in our study, we show that while newer models tend to generate more diverse claims, all models are less epistemically diverse than a basic web search. We find that model size has a negative impact on epistemic diversity, while retrieval-augmented generation (RAG) has a positive impact, though the improvement from RAG varies by the cultural context. Finally, compared to a traditional knowledge source (Wikipedia), we find that country-specific claims reflect the English language more than the local one, highlighting a gap in epistemic representation. [^1] Code and data: https://github.com/dwright37/llm-knowledge
@@ 1 Introduction
Figure 1: In this work, we measure epistemic diversity – via variability in claims about the world – for characterizing the risk of knowledge collapse with LLMs. This is done by comparing diversity across models, versions, settings, and knowledge sources.
@@
At the same time, recent studies have noted that LLM outputs are homogeneous Sourati et al. (2025b). For example, LLMs reflect only a narrow range of writing and reasoning styles, use a limited vocabulary Sourati et al. (2025a) and convey only certain semantics Lee et al. (2025); Jiang et al. (2025); Yu et al. (2023). This phenomenon, where models regress to a “central tendency”, may also affect knowledge i.e., the pieces of information that models tend to generate Dutta and Chakraborty (2023); Peterson (2025). If most people are exposed to largely the same information through LLMs, limiting that set of informaiton may diwinde our collective knowledge— a phenomenon broadly defined as knowledge collapse (Fig. 1). [^2] For an in depth definition and discussion, see Peterson (2025). Recent work has begun to theorize this Peterson (2025), and some empirical work indicates low knowledge diversity among LLMs in closed generation settings (e.g., recommendations of people and places Dudy et al. 2025; Poulain et al. 2024; Peterson 2025). However, there is no empirical work on LLMs’ open, unconstrained response generation where individual pieces of knowledge are extracted from the multiple claims which compose an LLM’s response. Additionally, there is no work characterizing how the diversity of this knowledge within LLMs has changed over time and across cultures. Such work could inform us about the current and future risk of knowledge collapse based on realistic user interactions (e.g., different people asking for information about the same topics in an open-ended fashion), as well as offer insights into how to prevent it.
@@
To address this, we perform an empirical study measuring the epistemic diversity of LLMs, defined as the diversity of claims about the world in a set of LLM outputs. As knowledge collapse is about collective knowledge, we simulate the scenario where different people prompt LLMs in different ways about the same topics and receive open-ended free-text responses embedded with multiple claims. As such, epistemic diversity is measured by (1) sampling responses from LLMs about a given topic using a set of 200 human-written prompt templates collected from Röttger et al. (2025), (2) decomposing individual LLM responses into one or more claims, (3) partitioning those claims into semantically equivalent clusters Farquhar et al. (2024), and (4) quantifying the diversity of the sample of claims with Hill-Shannon diversity, a widely used metric for measuring species diversity in ecology Roswell et al. (2021). To assess the risk of knowledge collapse, we use this methodology to study epistemic diversity in 27 LLMs spanning multiple versions, sizes, and release dates, and querying them for 155 topics. In doing so, we produce a dataset of 70M claims clustered by semantic equivalence generated by a set of the most popular LLMs from the past three years.
@@
For the selected topics, while we encouragingly find that epistemic diversity has increased since 2023 for three of four model families, it remains low for all models compared to the epistemic diversity of a rudimentary web search. Similar to contemporary work Zhang et al. (2025); Rassin et al. (2024), we find that model size has a statistically significant negative impact on epistemic diversity – smaller models generate more diverse knowledge than larger ones. In contrast, retrieval-augmented generation (RAG) has a statistically significant positive impact, highlighting the importance of RAG in preventing knowledge collapse. For country-specific topics, we find that RAG has an uneven effect; certain countries (e.g., the USA) see more benefit due to a greater diversity in their RAG sources. Finally, compared to a traditional knowledge source (Wikipedia) in both English and local languages for country-specific topics, we find that the claims generated in our study reflect English language knowledge more than local language knowledge, highlighting a gap in epistemic representation. In sum, for our topics, we find that there is a risk of knowledge collapse with current models given their low diversity and lack of epistemic representation, but that trends are improving and there are actionable ways to continue this improvement.
@@ 2 Background
An increasing body of work shows that LLMs suffer from a lack of diversity along many dimensions; for recent survey, see Sourati et al. (2025b), and recent benchmarking, see Guo et al. (2025). This lack of diversity includes lexical and stylistic Sourati et al. (2025a); Shaib et al. (2024); Padmakumar and He (2024), semantic Lee et al. (2025); Jiang et al. (2025); Yu et al. (2023); Padmakumar and He (2024); Moon et al. (2025), creative Xu et al. (2025); Moon et al. (2025); Wenger and Kenett (2025), conceptual (e.g., word-color associations) Murthy et al. (2025), recommendations Dudy et al. (2025); Poulain et al. (2024); Barolo et al. (2025), coding Shypula et al. (2025), and perspective diversity Wright et al. (2024); Abdurahman et al. (2024); Zhang et al. (2025); Durmus et al. (2023); Röttger et al. (2024); Moore et al. (2024); Hayati et al. (2024). Our work is most similar to perspective diversity, which has shown that LLM-generated views, opinions, and beliefs tend to reflect only a small subset of the world Durmus et al. (2023); Atari et al. ; Abdurahman et al. (2024); Alvero et al. (2024). Perspective diversity is usually measured with multiple-choice survey responses Durmus et al. (2023) or free-text responses partitioned into different categories Zhang et al. (2025).
Our work builds on this scholarship by looking at the diversity of all semantically equivalent classes of claims in open-ended free-text LLM responses. We avoid shallow and/or fuzzy features used in semantic and lexical similarity, and improve upon unreliable multiple-choice setups Röttger et al. (2024). Our methodology and empirical results show how this diversity has changed over time and across cultures, with actionable recommendations, and offering more general conclusions on knowledge diversity than previous work on closed-ended knowledge, semantic diversity, and perspective diversity.
@@
LLMs can exacerbate their own biases, leading to model collapse Shumailov et al. (2024); Guo et al. (2024). A growing concern among scholars is that such homogenization, combined with increased adoption of LLMs, will lead to epistemic problems at a societal level Zheng and Lee (2023); Messeri and Crockett (2024); Peterson (2025); Wagner and Jiang (2025); Qiu et al. (2025); Farrell et al. (2025). Peterson (2025) defines “knowledge collapse” as LLMs facilitating a dwindling of knowledge into an increasingly narrow set of ideas. Knowledge collapse may affect existing knowledge sources such as Wikipedia Wagner and Jiang (2025), erase minoritized knowledge Zheng and Lee (2023), pollute scientific discoveries Messeri and Crockett (2024), hamper political discourse Coeckelbergh (2025), and limit ideation in writing Anderson et al. (2024). These concerns echo other contemporary epistemic issues such as “popularity bias” Ciampaglia et al. (2018), and “filter bubbles” Nguyen et al. (2014) occurring with the use of recommender systems.
Our work assesses the risk of knowledge collapse with LLMs through a new methodology for measuring epistemic diversity and an empirical study of this phenomenon in LLMs. Using this, we compare previous prevailing means of knowledge acquisition and different cultures’ representation.
@@ 3 Problem Setup and Notation
From sports to math, people commonly query LLMs about similar topics. The responses that people receive are composed of claims about the topics, constituting individual pieces of information that people acquire from the LLMs. We measure the epistemic diversity of LLMs as the variation in these claims across different topics and prompts. To do so, we develop a methodology that reflects the common use case just outlined: we calculate the prevalence of sets of semantically equivalent claims among a broad set of open-ended, topic-specific LLM responses and measure their diversity in a statistically appropriate manner. Concretely, we transform these prevalences into categorical distributions Farquhar et al. (2024) and use Hill-Shannon Diversity, a statistically grounded measure of diversity used widely in ecology for measuring species diversity Peterson (2025); Roswell et al. (2021), to quantify the diversity of claims based on those distributions.
To describe our approach (Fig. 1), we adopt the following notation. First, a corpus of text $P_{m{}_{t{}}}{}$ is elicited about a topic $t$ from a model mm{}. $P_{m{}_{t{}}}{}$ contains free text which can be decomposed into a list of $n$ claims $C_{m{}_{t}}$, which can be further partitioned into a set of unique meaning classes $\mathcal{X}_{m{}_{t}}$. A unique meaning class is a cluster where all claims within a given class are semantically equivalent, and are not semantically equivalent to claims in other classes. Then, $x_{i}\in\mathcal{X}_{m_{t}}$ is defined as the empirical frequency of meaning class $i$ calculated from $C_{m{}_{t}}$, and $p_{i}$ is the probability of $i$, calculated as the relative frequency of $i$ in the sample, i.e., $\frac{x_{i}}{n}$.
Why semantic equivalence?
Our aim, reflected in our problem setup and formulation, is to measure how diverse are the unique pieces of knowledge that different people can be exposed to when interacting with an LLM. To contrast our methodology with similar contemporary work on diversity Lee et al. (2025); Wright et al. (2024); Jiang et al. (2025), an alternative version of our setup would be to partition claims based solely on semantic similarity, i.e., the distance between claims in a semantic embedding space. However, semantic similarity can only describe the degree to which two pieces of text are functionally similar, and only in extreme cases whether they convey the same piece of information. To illustrate: the phrases “Claude Shannon is the father of information theory” and “Claude Shannon is not the father of information theory” have high semantic similarity (0.94) [^3] Huggingface ID: all-MiniLM-L6-v2, even though they present opposing claims. Similarly, “Claude Shannon is the father of quantum theory” is an entirely different claim from “Claude Shannon is the father of information theory,” but these two also have high semantic similarity (0.805). Therefore, while semantic similarity can be useful in order to find candidate pairs which are semantically equivalent, partitioning based on similarity alone would lead to classes which contain potentially many different pieces of knowledge, thus underestimating diversity and not reflecting our definition of meaning classes. [^4] We will demonstrate this empirically by comparing our methodology to two previous approaches based solely on semantic similarity (Artificial Hivemind by Jiang et al. 2025 and LLM tropes by Wright et al. 2024).
@@ 4 Data Collection
    Cluster: Group the list of atomic claims into semantically equivalent meaning classes $\mathcal{X}_{m{}_{t}}$.
@@ 4.1 Generation and Decomposition
We use the open-ended writing assistance prompt templates collected in Röttger et al. (2025). The original prompt templates are sourced from WildChat Zhao et al. (2024), a collection of natural conversations between users and ChatGPT. From the original 1,000 writing assistance prompts, we manually select a subset of 479 prompts that primarily focus on information seeking and informational writing. We manually filter out not safe for work (NSFW) templates, creative writing templates such as “write a 50’s soviet style song about tt,” outlines such as “write an index for a book on tt,” or those explicitly asking for references. We then randomly sample 200 of these prompt templates, which are used to generate responses by inserting topic names into each template (every template is used for every topic in the study).
After generating responses ($P_{m{}_{t{}}}{}$), we decompose each response into a list of individual claims $C_{m{}_{t}}$. We do so using a strong open-weight LLM, [^5] Huggingface ID: meta-llama/Llama-3.1-70B-Instruct prompting the model to decompose non-overlapping chunks of three sentences at a time. This allows us to balance claim recall while ensuring that each input chunk is decontextualized.
@@
To evaluate the quality of decomposition, we develop three initial decomposition prompts (P1-P3) and have two independent annotators label (1) the quality of individual decomposed claims on a Likert scale from 1-5, and (2) how many claims in the original input chunk are missing from the list of decomposed claims. (On our scale, 1 means that the decomposed claim is not inferable from the original text, and 5 means that the claim is fully inferable and decontextualized.) The full annotation instructions appear in the Appendix § A.1. We label 100 input chunks for quality (1) and 108 instances for missing (2), all anonymized for the prompt type (P1-P3). Between the two annotators, we achieve a Kendall tau-b correlation Kendall (1938) of 0.53 for (1) and 0.39 for (2), indicating moderate to strong agreement. [^6] We select tau-b because it is suited for ordinal data with ranking ties Kendall (1945). We break ties by having a third annotator label those instances where the original annotators disagreed.
We then set up an LLM-as-a-judge using G-Eval Liu et al. (2023), which achieves 0.6 Pearson correlation with the human labels for decomposition quality and 0.68 for the number of missing claims. We use this to automatically label 6k instances across three decomposition prompt variants (P1-P3; see Appendix § A.2 for prompt text). The results are given in Table 1. We use P3 for the final decomposition in order to prioritize quality.
@@ 4.2 Clustering
Algorithm 1 outlines our approach for clustering the decomposed claims $C_{m{}_{t}}$ into semantically equivalent meaning classes $\mathcal{X}_{m{}_{t}}$
@@
Following Farquhar et al. (2024), we create clusters based on mutual entailment using a strong pretrained model for natural language inference. [^7] Huggingface ID: microsoft/deberta-large-mnli This is done by assuming that when a given claim mutually entails at least one claim in an existing cluster, it also entails all other claims in the cluster. To reduce the computational burden of measuring mutual entailment across $2n^{2}$ pairs for a set of $n$ claims, we only check entailment between the $N$ most similar claims (i.e., 2Nn2N*n comparisons), where similarity is measured using a strong S-BERT model. [^8] Huggingface ID: all-MiniLM-L6-v2 To mitigate potential drift in cluster cohesion (i.e., large clusters containing multiple meaning classes), we perform a post-processing step to break up large clusters using DBSCAN Ester et al. (1996), a related clustering approach Peterson (2025); Wright et al. (2024).
@@
Two independent annotators label: (1) the degree of cohesion among the clustered claims (Likert score between 1 and 5), and (2) whether any claims are missing from a given cluster (binary yes/no). We define cohesion as the degree to which each piece of text in the cluster conveys at least one piece of information in common. We define missing whether semantically similar claims which were not predicted to be included in a given cluster (i.e., singletons) should have been grouped with said cluster. See Appendix § A.1 for full instructions. We label 100 instances of cohesion (1) and 360 instances of missing (2), achieving a Kendall tau-b correlation of 0.38 for (1) and 87% agreement accuracy for (2). We again break ties using a third annotator and set up an LLM-as-a-judge using G-Eval in order to annotate and evaluate the rate of cluster cohesion and missing sentences in our final clusters. We develop prompts that achieve 0.68 Pearson $r$ with the human labels for measuring cluster cohesion and 83.5 weighted F1 score for measuring binary missing claims. 15,000 samples of 3-5 sentences are then automatically labeled from clusters acquired using Algorithm 1, where semantic similarity for task (2) is measured using S-BERT. We acquire a cluster cohesion score of 4.17 and a missing rate of 13.79%, indicating that the cluster cohesion is generally of high quality while missing only a relatively small proportion of singletons. We compare this to a semantic similarity only approach using DBSCAN in Appendix § A.4.1 Wright et al. (2024), showing that cluster cohesion degrades to 3.48, though with a lower missing rate of 0.4%.